### v1

# Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding Zusammenfassung

Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding
Zusammenfassung
keywords:
Introduction
Gold is a key metal in microelectronics due to its low resistivity, resistance to electromigration and chemical inertness. These properties make it an excellent choice for stable and reliable electrical contacts even under challenging environmental conditions 1. Chemical vapour deposition (CVD) is a technologically significant option for fabricating gold contacts or other metallic functional structures. CVD utilises the thermal decomposition of precursor molecules in a gaseous state 2. Molecules reaching the surface are thermally decomposed, leaving the central metal atom on the surface. If the decomposition is favored kinetically or thermodynamically on certain surfaces compared to others, the ensuing variation in reaction rates can be exploited for area selective deposition 2. This has been shown for various metals, e.g. Au on W 3, Au on Cr 4, or Cu on Pt 5. All of the aforementioned approaches require sophisticated lithography methods to define the growth surfaces, which can restrict the selection of substrate materials, ultimately constraining the degrees of freedom in device design 3, 6. Consequently, an area selective deposition technique that operates without the need for lithography, and is independent of the substrate type, would be highly advantageous. One option is to provide a localized energy input for CVD, which is frequently accomplished through laser irradiation 7, 8. However, the spatial resolution achievable through this method is typically restricted by the wavelength of the laser light employed. Although there has been recent evidence that this issue can be addressed in thermally decoupled 3D gold structures 9, the range of achievable structure geometries remains restricted. This is also applicable to local energy input through resistive heating 10. An alternative approach uses beams of charged particles, often electrons 11, 12. Instead of supplying additional energy during the CVD process, electrons can also be used to activate a specific surface 13, 14 or to deposit a seed layer for preferential decomposition of precursor molecules either by CVD 12, 15 or by atomic layer deposition (ALD) 16, 17. Seeding using a focused electron beam possesses significant potential, as direct electron beam writing is capable of operating on any surface topography that is even weakly conductive 18. Additionally, it provides the possibility for unparalleled spatial resolution 19. If the seeds previously deposited exhibit autocatalytic activity 12, 15, 20, 21, the temperature necessary for decomposition decreases considerably. This technique efficiently limits the CVD metal growth to a specific area, while also reducing the thermal load on the substrate. However, there is currently no proof of electron beam seeding resulting in autocatalytic growth for gold.
Results and Discussion
The principle of the proposed process is depicted in Figure 1. A seed layer is deposited by the impact of a focused electron beam on a sample surface at room or elevated temperature. The electron beam is then switched off while the molecular gas flow is maintained for a period of time. Subsequent investigations show autocatalytic growth of pure gold at both seeding temperatures.Electron Beam Seeding
Autocatalytic Growth of Gold
In order to achieve autocatalytic activity of the seed layer, two conditions must be met. Firstly, the previously deposited metallic nuclei must have a clean surface to participate in the reaction. Therefore, ultra-high vacuum (UHV) conditions are commonly used for this type of experiment. Additionally, all final reaction products that result from ligand cleavage must be volatile enough to desorb intactly 31. Secondly, the growth kinetics at the nucleus must promote chemisorption of the precursor molecules through the removal of the initial ligand with activation energy lower than the standard bond dissociation energies 15. This condition applies to some carbonyls, e.g. area selective CVD by electron beam seeding has been demonstrated for the deposition of transition metals using precursors like Fe(CO) 12, 21, Co(CO)NO 20, and Cr(CO) 15. In some cases high vacuum (HV) conditions were sufficient to observe autocatalysis, e.g. for Fe(CO) 40, but in others clean metal surfaces were additionally required excluding electron beam writing for seeding, e.g. for Co(CO) 41. Under ultra high vacuum conditions, NH dosing was found to inhibit autocatalytic growth, which could subsequently be reactivated by electron beam seeding, providing an additional control option in area selective CVD utilising Fe(CO) 21. Generally, carbonyls possess the thermodynamically stable CO ligand, making them the preferable choice for autocatalytically enhanced growth processes. In contrast, organic ligands such as the frequently used -diketonates (here acac) are not expected to have such favorable desorption properties as mentioned above. The autocatalytic growth experiments were performed in the same vacuum chamber. After depositing a square array for one temperature, the electron beam was switched off and the GIS nozzle was opened again for 30 min to deliver precursor molecules. Scanning electron (SE) micrographs for investigation were taken only after cooling the stage to room temperature and recovering the background pressure to the base pressure of about mbar. The experimental matrix continued with the next array, covering temperatures ranging from 30 °C up to 90 °C at intervals of 10 K. The growth results are shown in Figure 2(b) where carbon appears as a dark contrast and gold as a bright contrast, which is a typical characteristic of SE imaging. At 60 °C, no observable difference was noted when compared to the seed arrays deposited initially. However, at 70 °C, initial signs of CVD growth became evident. The contrast shifted from dark (carbon-rich) to bright (gold-rich), becoming more pronounced as the growth temperature increased. Squares with a greater number of repeats in the electron beam seeding are brighter, corresponding to more pronounced gold growth. This can be attributed to an increased density of gold nuclei available with deposition time. The optimal balance between growth rate and selectivity is achieved at a temperature of 80 °C. Raising the temperature further enhances the growth rate, but decreases selectivity due to the initiation of growth in the halo regions around the actual patterned squares. This phenomenon is demonstrated in Figure 2(c), a high-resolution SE micrograph of the square seeded at 90 °C with 20 repetitions at 5 nm pitch. Although there are no primary electrons in the halo, there are still secondary and backscattered electrons present with a significantly lower flux than in the deposition region 38. Consequently, numerous repetitions in seeding at 80 °C can also diminish the precision of square edges. Interestingly, electron beam seeding also works at room temperature. However, as anticipated, the higher desorption rate of ligands during seeding at elevated stage temperature suppresses carbon co-deposition, resulting in higher densities of catalytically active seeds. This is illustrated in Figures 1(a) and (b). Therefore, the films grown from seed layers deposited at room temperature show a significant carbon contrast, while at the elevated stage temperature the seed density is already high enough to form a quasi-continuous gold film, cf. Figure 2(b) middle row. For the tested pitch values and numbers of repeats, the growth of the autocatalytic film at a certain temperature is only influenced by the total electron dose. When comparing SE contrast and topography, it is evident that the seeds, which were deposited with different pitches and repeats but the same total dose, display the same thickness and SE contrast (grain size) after growth. The optimum dose is reached at about 140 pC/µm or 875 electrons per nm. Additionally, the process exhibits exceptional selectivity as gold growth is restricted to the presence of gold nuclei under these temperatures. It is worth noting, however, that a substantial decrease in CVD growth rate was observed over time, resulting in saturation for attainable thicknesses at approximately 6.5 nm. This could be due to surface poisoning, i.e. carbon coverage of the active gold nuclei due to residual hydrocarbons in the vacuum chamber and/or low desorption rates of the ligand fragments. In this instance, an improvement in the background vacuum would be promising. Another potential reason is the nuclei changing in size during autocatalytic growth. Size effects can have a significant impact on the catalytic activity of metallic nanoparticles, especially in the case of gold 42. It is worth mentioning that the observed decreasing CVD growth rate is not an obstacle to technological implementation, as seeding and growth can be easily iterated several times (as shown in Figure SI.3).
Beam-induced Heating and Autocatalysis
One of the main benefits of using a tightly focused electron beam to decompose molecules from the gas phase is its extreme 3D flexibility, resolution and shape fidelity 32. Generally, FEBID is considered to be stimulated by secondary electrons which dissociate the physisorbed precursor molecules. Electron-induced chemistry involves various reaction pathways, which depend on the local availability of electrons, molecules, and co-reactants 25, 31. Beam-induced heating is also a significant factor. As the growth surface locally heats up, the dynamics of the supplied precursor molecules and fragments will shift, e.g. towards faster diffusion and higher desorption rates. Estimates show that even slight changes in surface temperature, around 10 K, can significantly impact the vertical growth rate during 3D FEBID 43, 44, 45, 46, which typically operates in the reaction-rate limited growth regime. However, there is no experimental data available regarding the actual local temperature changes that may occur. Here, we exploit the onset of autocatalytic CVD during the growth of one-dimensional vertical nanostructures (pillars) under steady-state growth conditions as a means of in situ temperature measurement.
Conclusions
In this work, area selective CVD of gold by electron beam seeding is demonstrated for the first time. Gold nuclei, acting as a seed layer, are deposited through a maskless procedure by means of a focused electron beam, followed by gentle heating of the substrate, which triggers autocatalytic gold growth only in the seeded regions. The use of a focused electron beam ensures unparalleled spatial resolution, unhampered by mask usage typical in electron beam lithography and rendering complex multi-stage lithographic processing unnecessary. Seeding and subsequent growth both use the same precursor compound, Au(acac)Me, which is commercially available, long-term stable and non-toxic. The technique has been demonstrated to operate at low temperatures of approximately 80 °C, more than 200 K below the typical substrate temperature used in the established CVD process for this precursor. Moreover, it demands solely a high vacuum, making technological implementation remarkably straightforward. Any commercially available scanning electron microscope can be used for this method if it is equipped with a gas injection system and a heating stage. In the future, this process could be seamlessly integrated into conventional CVD reactors by equipping them with an electron gun featuring programmable position control. Unlike many other area selective CVD processes, direct electron beam writing for seeding is not affected by the surface material or topography. To demonstrate this capability, the same precursor was used to fabricate vertical nanostructures with high aspect ratios using the focused electron beam. In addition to achieving autocatalytically active seeds in 3D, the first quantitative measurement of the temperature change during direct electron beam writing of 3D structures was performed. This quantitative determination of the up to now inaccessible local temperature evolution is a major breakthrough as it enables improved modelling for accurate shape 3D nanoprinting through FEBID within the additive manufacturing field at a sub-100 nm scale 47, 49, 44, 45, 46. To summarise, this work presents a new, technologically powerful and easy to implement approach to area selective gold deposition. The low temperature necessary makes it highly suitable for processing innovative flexible devices that rely on heat-sensitive polymers, including flexible electronics 50 and flexible solar cells 51. Another exciting area of application is utilizing the maskless direct write seeding on 3D architectures, which enables the fabrication of nanostructures or localized electric contacts on intricate 3D devices.Acknowledgements
This work was funded by the German Research Association under grant agreement no. HO 5461/3-1 and by the Swiss National Fund by the COST-SNF project IZCOZ0_205450. We want to furthermore acknowledge support by the EU COST action CA 19140 ’FIT4NANO’ (www-fit4nano.eu). The authors are particularly grateful to Caroline Hain for her support with the English language.Conflict of Interest
Authors declare no conflict of interest. Keywords:Literatur
- [1] T. H. Baum, P. B. Comita, Chemical Vapor Deposition of Gold and Silver, chapter 6, pages 303–327, John Wiley & Sons, Ltd 1994.
- [2] T. T. Kodas, M. J. Hampden-Smith, Overview of Metal CVD, chapter 9, pages 429–498, John Wiley & Sons, Ltd 1994.
- [3] S. O. Colgate, G. J. Palenik, V. E. House, D. W. Schoenfeld, Journal of Vacuum Science & Technology A 1990, 8, 1411.
- [4] M. M. Banaszak Holl, P. F. Seidler, S. P. Kowalczyk, F. R. McFeely, Applied Physics Letters 1993, 62, 1475.
- [5] B. Lecohier, B. Calpini, J. Philippoz, H. van den Bergh, Journal of Applied Physics 1992, 72, 2022.
- [6] W. L. Gladfelter, Chemistry of Materials 1993, 5, 1372.
- [7] T. H. Baum, C. R. Jones, Applied Physics Letters 1985, 47, 538.
- [8] S. J. C. Irvine, D. Lamb, Photo-assisted CVD, in Chemical Vapour Deposition: Precursors, Processes and Applications, The Royal Society of Chemistry 2008.
- [9] J. Lasseter, P. D. Rack, S. J. Randolph, Nanomaterials 2023, 13.
- [10] F. Porrati, S. Barth, G. C. Gazzadi, S. Frabboni, O. M. Volkov, D. Makarov, M. Huth, ACS Nano 2023, 17, 4704, pMID: 36826847.
- [11] R. R. Kunz, T. E. Allen, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1987, 5, 1427.
- [12] R. R. Kunz, T. M. Mayer, Applied Physics Letters 1987, 50, 962.
- [13] M.-M. Walz, M. Schirmer, F. Vollnhals, T. Lukasczyk, H.-P. Steinrück, H. Marbach, Angewandte Chemie International Edition 2010, 49, 4669.
- [14] K. Muthukumar, H. O. Jeschke, R. Valentí, E. Begun, J. Schwenk, F. Porrati, M. Huth, Beilstein Journal of Nanotechnology 2012, 3, 546.
- [15] R. R. Kunz, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1988, 6, 1557.
- [16] A. J. M. Mackus, J. J. L. Mulders, M. C. M. van de Sanden, W. M. M. Kessels, Journal of Applied Physics 2010, 107, 116102.
- [17] A. Mameli, B. Karasulu, M. A. Verheijen, B. Barcones, B. Macco, A. J. M. Mackus, W. M. M. E. Kessels, F. Roozeboom, Chemistry of Materials 2019, 31, 1250.
- [18] I. Utke, P. Hoffmann, J. Melngailis, J. Vac. Sci. Technol. B 2008, 26, 1197.
- [19] W. F. van Dorp, C. W. Hagen, P. A. Crozier, P. Kruit, Nanotechnology 2008, 19, 225305.
- [20] F. Vollnhals, M. Drost, F. Tu, E. Carrasco, A. Späth, R. H. Fink, H.-P. Steinrück, H. Marbach, Beilstein Journal of Nanotechnology 2014, 5, 1175.
- [21] P. Martinović, L. Barnewitz, M. Rohdenburg, P. Swiderek, Journal of Vacuum Science & Technology A 2023, 41, 033207.
- [22] E. Feurer, H. Suhr, Applied Physics A 1987, 44, 171.
- [23] C. E. Larson, T. H. Baum, R. L. Jackson, MRS Online Proceedings Library (OPL) 1986, 75, 721.
- [24] A. Botman, J. J. L. Mulders, C. W. Hagen, Nanotechnology 2009, 20, 17pp.
- [25] S. Barth, M. Huth, F. Jungwirth, J. Mater. Chem. C 2020, 8, 15884.
- [26] J. J. L. Mulders, J. M. Veerhoek, E. G. T. Bosch, P. H. F. Trompenaars, Journal of Physics D: Applied Physics 2012, 45, 475301.
- [27] I. Utke, P. Hoffmann, B. Dwir, K. Leifer, E. Kapon, P. Doppelt, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures 2000, 18, 3168.
- [28] A. Marashdeh, T. Tiesma, N. J. C. van Velzen, S. Harder, R. W. A. Havenith, J. T. M. De Hosson, W. F. van Dorp, Beilstein Journal of Nanotechnology 2017, 8, 2753.
- [29] D. McIntosh, G. A. Ozin, Inorganic Chemistry 1977, 16, 51.
- [30] L. Jiang, Q. Xu, The Journal of Physical Chemistry A 2005, 109, 1026.
- [31] I. Utke, P. Swiderek, K. Höflich, K. Madajska, J. Jurczyk, P. Martinović, I. Szymańska, Coordination Chemistry Reviews 2022, 458, 213851.
- [32] R. Winkler, J. D. Fowlkes, P. D. Rack, H. Plank, Journal of Applied Physics 2019, 125, 210901.
- [33] O. Sqalli, I. Utke, P. Hoffmann, F. Marquis-Weible, Journal of Applied Physics 2002, 92, 1078.
- [34] J. J. L. Mulders, L. M. Belova, A. Riazanova, Nanotechnology 2010, 22, 055302.
- [35] K. Höflich, R. B. Yang, A. Berger, G. Leuchs, S. Christiansen, Adv. Mater. 2011, 23, 2657.
- [36] R. Winkler, F. P. Schmidt, U. Haselmann, J. D. Fowlkes, B. B. Lewis, G. Kothleitner, P. D. Rack, H. Plank, ACS Applied Materials and Interfaces 2017, 9, 8233.
- [37] D. Kuhness, A. Gruber, R. Winkler, J. Sattelkow, H. Fitzek, I. Letofsky-Papst, G. Kothleitner, H. Plank, ACS Applied Materials and Interfaces 2021, 13, 1178.
- [38] K. Höflich, J. Jurczyk, M. Puydinger, Y. Zhang, C. Kapusta, I. Utke, M. Puydinger dos Santos, M. Götz, C. Guerra-Nuñez, J. Best, C. Kapusta, I. Utke, ACS Appl. Mater. Interfaces 2017, 9, 24071.
- [39] P. Martinović, M. Rohdenburg, A. Butrymowicz, S. Sarigül, P. Huth, R. Denecke, I. B. Szymanska, P. Swiderek, Nanomaterials 2022.
- [40] G. Hochleitner, H. D. Wanzenboeck, E. Bertagnolli, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures Processing, Measurement, and Phenomena 2008, 26, 939.
- [41] R. Córdoba, J. Sesé, M. Ibarra, J. De Teresa, Applied Surface Science 2012, 263, 242.
- [42] M. Haruta, Catal. Today 1997, 36, 153 .
- [43] S. J. Randolph, J. D. Fowlkes, P. D. Rack, Journal of Applied Physics 2005, 97, 124312.
- [44] E. Mutunga, R. Winkler, P. D. Rack, H. Plank, J. D. Fowlkes, ACS Nano 2019, 13, 5198 .
- [45] L. Skoric, D. Sanz-Hernández, F. Meng, C. Donnelly, S. Merino-Aceituno, A. Fernández-Pacheco, Nano Letters 2020, 20, 184.
- [46] J. D. Fowlkes, R. Winkler, P. D. Rack, H. Plank, ACS Omega 2023, 8, 3148 .
- [47] J. D. Fowlkes, R. Winkler, B. B. Lewis, M. G. Stanford, H. Plank, P. D. Rack, ACS Nano 2016, 10, 6163.
- [48] J. Robertson, Mater. Sci. Eng. R 2002, 37, 129.
- [49] L. Keller, M. Huth, Beilstein Journal of Nanotechnology 2018, 9, 2581.
- [50] B. D. Gates, Science 2009, 323, 1566.
- [51] M. Pagliaro, R. Ciriminna, G. Palmisano, Chem. Sus. Chem 2008, 1, 880 .
- [52] K. Rykaczewski, W. B. White, A. G. Fedorov, Journal of Applied Physics 2007, 101, 054307.
- [53] D. C. Joy, Monte Carlo modeling for electron microscopy and microanalysis, volume 9, Oxford University Press 1995.
- [54] R. Browning, T. Z. Li, B. Chui, J. Ye, R. F. W. Pease, Z. Czyżewski, D. C. Joy, J. Appl. Phys. 1994, 76, 2016.
- [55] L. Reimer, Measurement Science and Technology 2000, 11, 1826.
Introduction
Gold is a key metal in microelectronics due to its low resistivity, resistance to electromigration and chemical inertness. These properties make it an excellent choice for stable and reliable electrical contacts even under challenging environmental conditions 1. Chemical vapour deposition (CVD) is a technologically significant option for fabricating gold contacts or other metallic functional structures. CVD utilises the thermal decomposition of precursor molecules in a gaseous state 2. Molecules reaching the surface are thermally decomposed, leaving the central metal atom on the surface. If the decomposition is favored kinetically or thermodynamically on certain surfaces compared to others, the ensuing variation in reaction rates can be exploited for area selective deposition 2. This has been shown for various metals, e.g. Au on W 3, Au on Cr 4, or Cu on Pt 5. All of the aforementioned approaches require sophisticated lithography methods to define the growth surfaces, which can restrict the selection of substrate materials, ultimately constraining the degrees of freedom in device design 3, 6. Consequently, an area selective deposition technique that operates without the need for lithography, and is independent of the substrate type, would be highly advantageous.
One option is to provide a localized energy input for CVD, which is frequently accomplished through laser irradiation 7, 8. However, the spatial resolution achievable through this method is typically restricted by the wavelength of the laser light employed. Although there has been recent evidence that this issue can be addressed in thermally decoupled 3D gold structures 9, the range of achievable structure geometries remains restricted. This is also applicable to local energy input through resistive heating 10. An alternative approach uses beams of charged particles, often electrons 11, 12.
Instead of supplying additional energy during the CVD process, electrons can also be used to activate a specific surface 13, 14 or to deposit a seed layer for preferential decomposition of precursor molecules either by CVD 12, 15 or by atomic layer deposition (ALD) 16, 17. Seeding using a focused electron beam possesses significant potential, as direct electron beam writing is capable of operating on any surface topography that is even weakly conductive 18. Additionally, it provides the possibility for unparalleled spatial resolution 19. If the seeds previously deposited exhibit autocatalytic activity 12, 15, 20, 21, the temperature necessary for decomposition decreases considerably. This technique efficiently limits the CVD metal growth to a specific area, while also reducing the thermal load on the substrate. However, there is currently no proof of electron beam seeding resulting in autocatalytic growth for gold.
In this study, area selective CVD of gold by electron beam seeding is demonstrated for the first time. The process employs a focused electron beam for the non-thermal electron-triggered decomposition of Au(acac)Me which results in autocatalytically active gold nuclei. Subsequently, the same precursor is thermally decomposed at the gold seeds. This autocatalytic CVD of gold is observed at relatively low substrate temperatures of about 80 °C and requires only high vacuum conditions. Electron beam seeding is highly versatile, as demonstrated by the deposition of high aspect ratio pillars using the same precursor. The temperature-dependent autocatalytic CVD in this case provides the first quantitative measurement of the temperature change induced by the electron beam during the direct writing of 3D structures.
Results and Discussion
The principle of the proposed process is depicted in Figure 1. A seed layer is deposited by the impact of a focused electron beam on a sample surface at room or elevated temperature. The electron beam is then switched off while the molecular gas flow is maintained for a period of time. Subsequent investigations show autocatalytic growth of pure gold at both seeding temperatures.
Electron Beam Seeding
All experiments were carried out in a conventional dual beam instrument fitted with a gas injection system to introduce precursor molecules locally and a sample heating stage, as detailed in the Supporting Information (SI). The autocatalytically enhanced CVD process was found to work reliably on all used bulk substrates. Nevertheless, it is crucial to clean the sample surface prior to the process. Any organic or inorganic residues on the surface hinder the reproducibility of the process by partially impeding autocatalysis, or by functioning as preferred precursor chemisorption sites and thus "contamination seeds."
The precursor compound used, Au(acac)Me; dimethyl(acetylacetonate)gold(III), is primarily known for its use in CVD where it provides pure gold deposition in plasma assisted processing at low temperatures 22. The thermal CVD process was carried out at a substrate temperature of 300 °C. It produced gold films of medium quality with 85 at.% gold content. However, the corresponding fluorinated diketonates, namely tfac and hfac (tetra- and hexafluoroacetonate), resulted in pure gold deposition 23. Au(acac)Me is also commonly utilised in focused electron beam induced deposition (FEBID), a direct writing technique where the precursor molecules physisorbed to the surface are dissociated using a focused electron beam 18.
As electron-induced dissociation is not bonding selective and occurs at room temperature, the use of metal-organic precursor compounds often results in a carbon-rich deposit 24. Efforts have been made to deposit materials with inorganic ligands, which have shown promise especially for magnetic materials 25. For gold, the inorganic precursor, Au(CO)Cl, containing both chlorine and carbonyl ligands, allows high-purity deposition using focused electron beam writing. However, the chlorine content restricts spatial selectivity due to etching effects 26. Similar to PFAuCl, which also provides high-purity gold deposition 27, this compound is not stable under vacuum conditions, severely limiting its practical usability 28. It would be of significant interest to explore gold carbonyls in this context. Although they can be synthesised, gold carbonyls exist only in specific laboratory environments and are therefore not suitable for typical FEBID or CVD experiments 29, 30. For this reason, precursor compounds with organic ligands were used in most of the reported FEBID experiments 31.
The key advantage of the direct electron beam writing process is that it provides extreme shape flexibility in three dimensions 32. To efficiently utilize these distinctive 3D capabilities, several techniques have been employed to attain pure deposits. Here, we will particularly focus on the influence of heat, either as beam-induced heating, especially for large beam currents and thin 3D geometries, or applied externally by sample heating at the various stages of processing. Heat input is viewed as a means of enhancing deposit purity as it promotes the desorption rates of cleaved ligands and contributes to ligand dissociation by thermal assistance 31. However, the acac ligand has a low desorption rate and a tendency to polymerize under electron beam impact. This limits the achievable gold contents to approximately 30 at.% when using thermal energy input 33, 34. To achieve pure 3D gold structures, additional oxidizing agents must be used in combination with elaborate shape optimization for effective purification 35, 36, 37. The key question is whether this presents a fundamental constraint, or whether thermal assistance at low electron fluxes, as in the case of silver deposition 38, 39, can result in pure yet still selective metal deposition. In this scenario, the clean surface of the gold seed could even be autocatalytically active.
To investigate this hypothesis, the following experimental matrix was performed. Under a continuous gas flow, an array of 2x2 µm squares was patterned with varying electron beam doses using a beam energy of 5 keV and a beam current of 100 pA. Figure 2(a) shows part of the array, the complete array is available in the SI. As a representation of the applied electron dose, the patterning time is provided in milliseconds within each square. The corresponding pixel spacing (pitch) and number of repetitions (repeats) are also indicated alongside. The utilized pitches of 10 and 5 nm correspond to doses of 6.25 and 25 electrons per nm and repeat, respectively. Starting from a single repeat which is supposed to deposit only isolated gold atoms and small clusters up to 100 repeats were conducted, as the latter would lead to a thin FEBID deposit. To investigate the effect of ligand desorption, the seed deposition was carried out at room temperature (30 °C) and at elevated stage temperatures. Similar to early reports on Fe(CO) by Kunz 12, we observed efficient electron-enhanced CVD through continued electron irradiation under low fluxes and elevated stage temperatures (see Figure SI.3). Therefore special care was taken to avoid unintended beam exposure, and no imaging was conducted before or after the seed array deposition. As the GIS system has an automatically controlled valve, gas flow occurred only during seeding for the array deposition time of 1 min 18 s. As a control experiment for electron beam induced surface activation, the same array was deposited at the different substrate temperatures without gas flow (see Figure SI.4).
Autocatalytic Growth of Gold
In order to achieve autocatalytic activity of the seed layer, two conditions must be met. Firstly, the previously deposited metallic nuclei must have a clean surface to participate in the reaction. Therefore, ultra-high vacuum (UHV) conditions are commonly used for this type of experiment. Additionally, all final reaction products that result from ligand cleavage must be volatile enough to desorb intactly 31. Secondly, the growth kinetics at the nucleus must promote chemisorption of the precursor molecules through the removal of the initial ligand with activation energy lower than the standard bond dissociation energies 15.
This condition applies to some carbonyls, e.g. area selective CVD by electron beam seeding has been demonstrated for the deposition of transition metals using precursors like Fe(CO) 12, 21, Co(CO)NO 20, and Cr(CO) 15. In some cases high vacuum (HV) conditions were sufficient to observe autocatalysis, e.g. for Fe(CO) 40, but in others clean metal surfaces were additionally required excluding electron beam writing for seeding, e.g. for Co(CO) 41. Under ultra high vacuum conditions, NH dosing was found to inhibit autocatalytic growth, which could subsequently be reactivated by electron beam seeding, providing an additional control option in area selective CVD utilising Fe(CO) 21. Generally, carbonyls possess the thermodynamically stable CO ligand, making them the preferable choice for autocatalytically enhanced growth processes. In contrast, organic ligands such as the frequently used -diketonates (here acac) are not expected to have such favorable desorption properties as mentioned above.
The autocatalytic growth experiments were performed in the same vacuum chamber. After depositing a square array for one temperature, the electron beam was switched off and the GIS nozzle was opened again for 30 min to deliver precursor molecules.
Scanning electron (SE) micrographs for investigation were taken only after cooling the stage to room temperature and recovering the background pressure to the base pressure of about mbar. The experimental matrix continued with the next array, covering temperatures ranging from 30 °C up to 90 °C at intervals of 10 K.
The growth results are shown in Figure 2(b) where carbon appears as a dark contrast and gold as a bright contrast, which is a typical characteristic of SE imaging. At 60 °C, no observable difference was noted when compared to the seed arrays deposited initially. However, at 70 °C, initial signs of CVD growth became evident. The contrast shifted from dark (carbon-rich) to bright (gold-rich), becoming more pronounced as the growth temperature increased. Squares with a greater number of repeats in the electron beam seeding are brighter, corresponding to more pronounced gold growth. This can be attributed to an increased density of gold nuclei available with deposition time. The optimal balance between growth rate and selectivity is achieved at a temperature of 80 °C. Raising the temperature further enhances the growth rate, but decreases selectivity due to the initiation of growth in the halo regions around the actual patterned squares. This phenomenon is demonstrated in Figure 2(c), a high-resolution SE micrograph of the square seeded at 90 °C with 20 repetitions at 5 nm pitch. Although there are no primary electrons in the halo, there are still secondary and backscattered electrons present with a significantly lower flux than in the deposition region 38. Consequently, numerous repetitions in seeding at 80 °C can also diminish the precision of square edges.
Interestingly, electron beam seeding also works at room temperature. However, as anticipated, the higher desorption rate of ligands during seeding at elevated stage temperature suppresses carbon co-deposition, resulting in higher densities of catalytically active seeds. This is illustrated in Figures 1(a) and (b). Therefore, the films grown from seed layers deposited at room temperature show a significant carbon contrast, while at the elevated stage temperature the seed density is already high enough to form a quasi-continuous gold film, cf. Figure 2(b) middle row.
For the tested pitch values and numbers of repeats, the growth of the autocatalytic film at a certain temperature is only influenced by the total electron dose. When comparing SE contrast and topography, it is evident that the seeds, which were deposited with different pitches and repeats but the same total dose, display the same thickness and SE contrast (grain size) after growth. The optimum dose is reached at about 140 pC/µm or 875 electrons per nm. Additionally, the process exhibits exceptional selectivity as gold growth is restricted to the presence of gold nuclei under these temperatures. It is worth noting, however, that a substantial decrease in CVD growth rate was observed over time, resulting in saturation for attainable thicknesses at approximately 6.5 nm. This could be due to surface poisoning, i.e. carbon coverage of the active gold nuclei due to residual hydrocarbons in the vacuum chamber and/or low desorption rates of the ligand fragments. In this instance, an improvement in the background vacuum would be promising. Another potential reason is the nuclei changing in size during autocatalytic growth. Size effects can have a significant impact on the catalytic activity of metallic nanoparticles, especially in the case of gold 42. It is worth mentioning that the observed decreasing CVD growth rate is not an obstacle to technological implementation, as seeding and growth can be easily iterated several times (as shown in Figure SI.3).
The quality of the gold films obtained was later assessed by scanning transmission electron microscopy (STEM). Figure 3 depicts cross-sectional images of films seeded at (a) 30 and (b) 90 °C, both grown at 90 °C. The high-angle annular dark-field (HAADF) images prove the presence of pure crystalline gold in both instances, recognizable as the periodic atomic order in bright contrast regions. The close-up displayed in Figure 3(c) of a distinct grain from the halo region permits the evaluation of the crystal structure through the acquisition of a fast-Fourier transform (FFT) image that can be indexed based on the lattice parameters of single-crystalline gold.
Beam-induced Heating and Autocatalysis
One of the main benefits of using a tightly focused electron beam to decompose molecules from the gas phase is its extreme 3D flexibility, resolution and shape fidelity 32. Generally, FEBID is considered to be stimulated by secondary electrons which dissociate the physisorbed precursor molecules. Electron-induced chemistry involves various reaction pathways, which depend on the local availability of electrons, molecules, and co-reactants 25, 31. Beam-induced heating is also a significant factor. As the growth surface locally heats up, the dynamics of the supplied precursor molecules and fragments will shift, e.g. towards faster diffusion and higher desorption rates. Estimates show that even slight changes in surface temperature, around 10 K, can significantly impact the vertical growth rate during 3D FEBID 43, 44, 45, 46, which typically operates in the reaction-rate limited growth regime. However, there is no experimental data available regarding the actual local temperature changes that may occur. Here, we exploit the onset of autocatalytic CVD during the growth of one-dimensional vertical nanostructures (pillars) under steady-state growth conditions as a means of in situ temperature measurement.
Figure 4(a)-(c) shows example images of deposited nanopillars using the precursor Au(acac)Me with diameters of about 50 nm and aspect ratios of 40:1 and larger. The primary energy used was 15 keV at a beam current of 100 pA, which aligns with typical standards in 3D FEBID printing. Stationary growth conditions were achieved using a dwell time of 26.2 ms 43, 47. Each pillar’s overall growth duration was consistently maintained at 2 min 30 s deposition time.
Figure 4 (a) presents the nanopillar obtained at room temperature, showing clear bright particles in the upper part. The presence of gold particles confirms the existence of autocatalytically active gold nuclei on the surface of the pillar and the attainment of the temperature threshold T for visible particle growth. By adjusting the substrate temperature, the region of particle growth onset can be moved along the pillar, as demonstrated by the white arrows in Figures 4(a)-(c). The corresponding substrate temperatures are subsequently plotted as a function of the vertical distance from the substrate surface to the onset region, cf. Figure 4(d). The resulting decrease in with the increase in substrate temperature is linear.
This linear relationship is in line with the fact that there is a negligible loss of thermal energy across the surface of the pillar (refer to SI for estimates on radiative and convective heat losses). Therefore, all the thermal energy introduced by the stationary electron beam is conducted through the pillar into the substrate, which can be considered as a constant temperature heat sink. The heating induced by the beam during growth results in a linear temperature gradient within the pillar 18. Therefore, the vertical position of the CVD onset region is dependent on the stage temperature in a linear manner. The overall height of the pillars reduces for higher stage temperatures, as a result of the increased desorption rate of precursor molecules.
The intersection point of the linear fit with the temperature axis yields a measured temperature of T 112 °C, which indicates the point at which particle growth becomes visible. The thermal decomposition of the precursor conforms to Arrhenius’ law, with the autocatalysis rate exponentially dependent on temperature , where the activation energy acts as the threshold value and is the Boltzmann constant. The overall growth duration was relatively short, requiring a significant growth rate to observe particle growth. These findings align with our observations of the electron beam-seeded autocatalytic growth at temperatures of 100-110 °C, which was highly efficient but showed decreasing selectivity. For 3D growth, reduced selectivity is not an issue since only the thermally decoupled areas reach high enough temperatures.
The proportionality constant K/nm indicates the temperature variation per unit length along the pillar. It offers a quantitative measure of the thermal energy introduced by the electron beam, providing valuable insights for modelling 3D electron beam printing and the possibility of quantitatively determining the thermal conductivity of the deposited material. For this purpose, the measurements are supported by Monte-Carlo modelling of the electron trajectories obtained for this geometry. A mass density range of gcm is assumed, corresponding to the typically observed Au contents in deposits of this precursor, cf. SI for more details on the Monte-Carlo modelling. The obtained range for the average inelastic energy loss keV per incident electron inside the nanostructure can then be used to determine the thermal conductivity of the deposit material:
with the electron beam current , elementary charge and the cross-sectional area of the nanopillar. This can be considered an upper limit since some inelastic losses are due to the production of secondary electrons (and X-rays), while the rest results in Joule heating (e.g. from an assessment of this geometry 44).
The value 0.4 W/(Km) obtained for thermal conductivity has an uncertainty margin of approximately 24 % and falls within the commonly observed range for various carbon modifications, particularly those ranging from amorphous to diamond-like carbon 48. By repeating this type of experiment for different acceleration voltages, beam currents and possibly other precursors, the thermal conductivity of the FEBID material for the different deposition conditions can be determined quantitatively.
Conclusions
In this work, area selective CVD of gold by electron beam seeding is demonstrated for the first time. Gold nuclei, acting as a seed layer, are deposited through a maskless procedure by means of a focused electron beam, followed by gentle heating of the substrate, which triggers autocatalytic gold growth only in the seeded regions. The use of a focused electron beam ensures unparalleled spatial resolution, unhampered by mask usage typical in electron beam lithography and rendering complex multi-stage lithographic processing unnecessary. Seeding and subsequent growth both use the same precursor compound, Au(acac)Me, which is commercially available, long-term stable and non-toxic. The technique has been demonstrated to operate at low temperatures of approximately 80 °C, more than 200 K below the typical substrate temperature used in the established CVD process for this precursor. Moreover, it demands solely a high vacuum, making technological implementation remarkably straightforward. Any commercially available scanning electron microscope can be used for this method if it is equipped with a gas injection system and a heating stage. In the future, this process could be seamlessly integrated into conventional CVD reactors by equipping them with an electron gun featuring programmable position control.
Unlike many other area selective CVD processes, direct electron beam writing for seeding is not affected by the surface material or topography. To demonstrate this capability, the same precursor was used to fabricate vertical nanostructures with high aspect ratios using the focused electron beam. In addition to achieving autocatalytically active seeds in 3D, the first quantitative measurement of the temperature change during direct electron beam writing of 3D structures was performed. This quantitative determination of the up to now inaccessible local temperature evolution is a major breakthrough as it enables improved modelling for accurate shape 3D nanoprinting through FEBID within the additive manufacturing field at a sub-100 nm scale 47, 49, 44, 45, 46.
To summarise, this work presents a new, technologically powerful and easy to implement approach to area selective gold deposition. The low temperature necessary makes it highly suitable for processing innovative flexible devices that rely on heat-sensitive polymers, including flexible electronics 50 and flexible solar cells 51. Another exciting area of application is utilizing the maskless direct write seeding on 3D architectures, which enables the fabrication of nanostructures or localized electric contacts on intricate 3D devices.
Acknowledgements
This work was funded by the German Research Association under grant agreement no. HO 5461/3-1 and by the Swiss National Fund by the COST-SNF project IZCOZ0_205450. We want to furthermore acknowledge support by the EU COST action CA 19140 ’FIT4NANO’ (www-fit4nano.eu). The authors are particularly grateful to Caroline Hain for her support with the English language.
Conflict of Interest
Authors declare no conflict of interest.
Keywords:
Literatur
- [1] T. H. Baum, P. B. Comita, Chemical Vapor Deposition of Gold and Silver, chapter 6, pages 303–327, John Wiley & Sons, Ltd 1994.
- [2] T. T. Kodas, M. J. Hampden-Smith, Overview of Metal CVD, chapter 9, pages 429–498, John Wiley & Sons, Ltd 1994.
- [3] S. O. Colgate, G. J. Palenik, V. E. House, D. W. Schoenfeld, Journal of Vacuum Science & Technology A 1990, 8, 1411.
- [4] M. M. Banaszak Holl, P. F. Seidler, S. P. Kowalczyk, F. R. McFeely, Applied Physics Letters 1993, 62, 1475.
- [5] B. Lecohier, B. Calpini, J. Philippoz, H. van den Bergh, Journal of Applied Physics 1992, 72, 2022.
- [6] W. L. Gladfelter, Chemistry of Materials 1993, 5, 1372.
- [7] T. H. Baum, C. R. Jones, Applied Physics Letters 1985, 47, 538.
- [8] S. J. C. Irvine, D. Lamb, Photo-assisted CVD, in Chemical Vapour Deposition: Precursors, Processes and Applications, The Royal Society of Chemistry 2008.
- [9] J. Lasseter, P. D. Rack, S. J. Randolph, Nanomaterials 2023, 13.
- [10] F. Porrati, S. Barth, G. C. Gazzadi, S. Frabboni, O. M. Volkov, D. Makarov, M. Huth, ACS Nano 2023, 17, 4704, pMID: 36826847.
- [11] R. R. Kunz, T. E. Allen, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1987, 5, 1427.
- [12] R. R. Kunz, T. M. Mayer, Applied Physics Letters 1987, 50, 962.
- [13] M.-M. Walz, M. Schirmer, F. Vollnhals, T. Lukasczyk, H.-P. Steinrück, H. Marbach, Angewandte Chemie International Edition 2010, 49, 4669.
- [14] K. Muthukumar, H. O. Jeschke, R. Valentí, E. Begun, J. Schwenk, F. Porrati, M. Huth, Beilstein Journal of Nanotechnology 2012, 3, 546.
- [15] R. R. Kunz, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1988, 6, 1557.
- [16] A. J. M. Mackus, J. J. L. Mulders, M. C. M. van de Sanden, W. M. M. Kessels, Journal of Applied Physics 2010, 107, 116102.
- [17] A. Mameli, B. Karasulu, M. A. Verheijen, B. Barcones, B. Macco, A. J. M. Mackus, W. M. M. E. Kessels, F. Roozeboom, Chemistry of Materials 2019, 31, 1250.
- [18] I. Utke, P. Hoffmann, J. Melngailis, J. Vac. Sci. Technol. B 2008, 26, 1197.
- [19] W. F. van Dorp, C. W. Hagen, P. A. Crozier, P. Kruit, Nanotechnology 2008, 19, 225305.
- [20] F. Vollnhals, M. Drost, F. Tu, E. Carrasco, A. Späth, R. H. Fink, H.-P. Steinrück, H. Marbach, Beilstein Journal of Nanotechnology 2014, 5, 1175.
- [21] P. Martinović, L. Barnewitz, M. Rohdenburg, P. Swiderek, Journal of Vacuum Science & Technology A 2023, 41, 033207.
- [22] E. Feurer, H. Suhr, Applied Physics A 1987, 44, 171.
- [23] C. E. Larson, T. H. Baum, R. L. Jackson, MRS Online Proceedings Library (OPL) 1986, 75, 721.
- [24] A. Botman, J. J. L. Mulders, C. W. Hagen, Nanotechnology 2009, 20, 17pp.
- [25] S. Barth, M. Huth, F. Jungwirth, J. Mater. Chem. C 2020, 8, 15884.
- [26] J. J. L. Mulders, J. M. Veerhoek, E. G. T. Bosch, P. H. F. Trompenaars, Journal of Physics D: Applied Physics 2012, 45, 475301.
- [27] I. Utke, P. Hoffmann, B. Dwir, K. Leifer, E. Kapon, P. Doppelt, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures 2000, 18, 3168.
- [28] A. Marashdeh, T. Tiesma, N. J. C. van Velzen, S. Harder, R. W. A. Havenith, J. T. M. De Hosson, W. F. van Dorp, Beilstein Journal of Nanotechnology 2017, 8, 2753.
- [29] D. McIntosh, G. A. Ozin, Inorganic Chemistry 1977, 16, 51.
- [30] L. Jiang, Q. Xu, The Journal of Physical Chemistry A 2005, 109, 1026.
- [31] I. Utke, P. Swiderek, K. Höflich, K. Madajska, J. Jurczyk, P. Martinović, I. Szymańska, Coordination Chemistry Reviews 2022, 458, 213851.
- [32] R. Winkler, J. D. Fowlkes, P. D. Rack, H. Plank, Journal of Applied Physics 2019, 125, 210901.
- [33] O. Sqalli, I. Utke, P. Hoffmann, F. Marquis-Weible, Journal of Applied Physics 2002, 92, 1078.
- [34] J. J. L. Mulders, L. M. Belova, A. Riazanova, Nanotechnology 2010, 22, 055302.
- [35] K. Höflich, R. B. Yang, A. Berger, G. Leuchs, S. Christiansen, Adv. Mater. 2011, 23, 2657.
- [36] R. Winkler, F. P. Schmidt, U. Haselmann, J. D. Fowlkes, B. B. Lewis, G. Kothleitner, P. D. Rack, H. Plank, ACS Applied Materials and Interfaces 2017, 9, 8233.
- [37] D. Kuhness, A. Gruber, R. Winkler, J. Sattelkow, H. Fitzek, I. Letofsky-Papst, G. Kothleitner, H. Plank, ACS Applied Materials and Interfaces 2021, 13, 1178.
- [38] K. Höflich, J. Jurczyk, M. Puydinger, Y. Zhang, C. Kapusta, I. Utke, M. Puydinger dos Santos, M. Götz, C. Guerra-Nuñez, J. Best, C. Kapusta, I. Utke, ACS Appl. Mater. Interfaces 2017, 9, 24071.
- [39] P. Martinović, M. Rohdenburg, A. Butrymowicz, S. Sarigül, P. Huth, R. Denecke, I. B. Szymanska, P. Swiderek, Nanomaterials 2022.
- [40] G. Hochleitner, H. D. Wanzenboeck, E. Bertagnolli, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures Processing, Measurement, and Phenomena 2008, 26, 939.
- [41] R. Córdoba, J. Sesé, M. Ibarra, J. De Teresa, Applied Surface Science 2012, 263, 242.
- [42] M. Haruta, Catal. Today 1997, 36, 153 .
- [43] S. J. Randolph, J. D. Fowlkes, P. D. Rack, Journal of Applied Physics 2005, 97, 124312.
- [44] E. Mutunga, R. Winkler, P. D. Rack, H. Plank, J. D. Fowlkes, ACS Nano 2019, 13, 5198 .
- [45] L. Skoric, D. Sanz-Hernández, F. Meng, C. Donnelly, S. Merino-Aceituno, A. Fernández-Pacheco, Nano Letters 2020, 20, 184.
- [46] J. D. Fowlkes, R. Winkler, P. D. Rack, H. Plank, ACS Omega 2023, 8, 3148 .
- [47] J. D. Fowlkes, R. Winkler, B. B. Lewis, M. G. Stanford, H. Plank, P. D. Rack, ACS Nano 2016, 10, 6163.
- [48] J. Robertson, Mater. Sci. Eng. R 2002, 37, 129.
- [49] L. Keller, M. Huth, Beilstein Journal of Nanotechnology 2018, 9, 2581.
- [50] B. D. Gates, Science 2009, 323, 1566.
- [51] M. Pagliaro, R. Ciriminna, G. Palmisano, Chem. Sus. Chem 2008, 1, 880 .
- [52] K. Rykaczewski, W. B. White, A. G. Fedorov, Journal of Applied Physics 2007, 101, 054307.
- [53] D. C. Joy, Monte Carlo modeling for electron microscopy and microanalysis, volume 9, Oxford University Press 1995.
- [54] R. Browning, T. Z. Li, B. Chui, J. Ye, R. F. W. Pease, Z. Czyżewski, D. C. Joy, J. Appl. Phys. 1994, 76, 2016.
- [55] L. Reimer, Measurement Science and Technology 2000, 11, 1826.
Supporting Information
Methods and Full Seed Array
Silicon wafers with native oxide and varied doping levels were employed as the substrate material and chipped into 1 cm x 1 cm pieces. As the doping level decides the conductivity and therefore the secondary electron yields, the minimum required electron doses for successful seeding varied. The silicon chips with a native oxide layer were vertically sonicated in N-Methyl-2-pyrrolidone for 20 min, followed by a rinse with isopropyl alcohol. The chips were then soaked in a 3:1 Piranha solution for at least 1 h before being rinsed with deionized water and dried using nitrogen gas. This process effectively removes any organic material from the surface of the sample.
All testing was carried out using a Helios 5 UX dual beam microscope. The gas injection system (GIS) with crystals of Au(acac)Me (CAS: 14951-50-9) was heated to 30 °C. To enable autocatalysis growth of gold films, gold nuclei were seeded with an electron beam at 5 kV acceleration voltage and a beam current of 100 pA. The base chamber pressure prior to opening the GIS valve was typically mbar, with the GIS valve open during autocatalysis - mbar at the start of the experiments, dropping to approximately at the end of an experimental day. The Kleindiek vacuum compatible Micro Heating and Cooling Stage (MHCS) was used together with a temperature control system to heat the sample up to 110 °C. The complete pattern array for all experiments is shown in Figure SI.1(a). It comprises 2 µm x 2 µm squares with variable pixel-to-pixel pitches (5, 10 and 20 nm) of the electron beam, separated by 2 µm, and different deposition times (achieved by scan repeats from 1 to 100). A non-linear color scale indicates the deposition time of a particular square, ranging from 10 ms to 16 s. The whole seed array takes 1 min 18 s of deposition time to complete. The assortment of squares within the full seed array enables the gold film thickness to be optimised and to understand how autocatalytic growth depends on the electron beam patterning. Squares featuring the same deposition time (and dose) are highlighted with a coloured frame.
The experimental layout of all autocatalytically grown films is shown schematically in Figure SI.1(b). The complete experimental procedure for seeding and autocatalytic growth at various temperatures is detailed below:
1.
Focus the beam onto a designated area with a small feature on it. It is crucial to refocus the beam before each seed deposition owing to material expansion caused by heat.
2.
Move the stage to the first patterning area, which should be at least 50 µm away from the focusing area (position i in Figure SI.1(b)), without conducting any observation with the SEM.
3.
Deposit the full seed array at room temperature, approximately 30 °C for all experiments.
4.
Increase the stage temperature to 40 °C.
5.
Allow autocatalytic growth of the gold films on the seeded areas with an open GIS and maintain a stable stage temperature for 30 min.
6.
Returning the stage to the focusing area, refocusing the beam, and capturing a SE image of the resulting films.
7.
Moving to the second patterning area, which is situated 100 µm from the previously grown films (position ii in Figure SI.1(b)).
8.
Deposition of the full seed array onto the hot stage (40 °C).
9.
Autocatalytic growth of the gold films on the seeded areas with open GIS and stable stage temperature for a duration of 30 min.
10.
Capturing SE images of the resulting films.
11.
Returning to the focusing area and replicating the procedure for elevated temperatures ranging from 50 to 90 °C at intervals of 10 K
Pillars for measuring the temperature induced by an electron beam were grown using a focused electron beam of 15 kV and 100 pA. Deposition time for all pillars was achieved by a spot deposition with 5725 repeats with the maximum possible dwell time of 26.2 ms, resulting in a total deposition time of about 2 min 30 s for a single pillar. The layout for experimental deposition of all pillars under different stage temperatures is illustrated in Figure SI.1(c). The entire experimental protocol for pillars growth is described below:
1.
Focus the beam on a designated area with a small feature. Refocus the beam before each seed deposition due to thermal material expansion.
2.
Without observation with the SEM, move the stage to the first patterning area, which is at least 20 µm away from the focusing area. Room temperature deposition is done at position i in Figure SI.1(c).
3.
Deposit three pillars with the same parameters.
4.
Close and retract the GIS, tilt the stage to 45°.
5.
Take high-resolution SEM images of all pillars.
6.
Tilt and move back to the focusing area.
7.
Repeat the same protocol for higher temperatures ranging from 40 to 110 °C with a step of 10 K.
8.
Ensure that all depositions are done in close proximity to the focusing spot to guarantee similar beam waist for each iteration.
SE Micrographs of Autocatalytically Grown Films on the Full Seed Arrays
SE micrographs of gold films, grown on the full seed arrays at various temperatures, are displayed in Figure SI.2. The gold films in the top row were grown on seed arrays deposited at a room temperature of 30 °C, while the seed arrays for the gold films in the bottom row were deposited at higher temperatures ranging from 60 to 90 °C. Autocatalytic growth was continued for 30 min for each film.
Time Evolution of Autocatalytic Growth on the Full Seed Arrays
The autocatalytic growth of gold films on two seed arrays for additional intermediate electron supply is shown in Figure SI.3, indicating the time evolution. The initial SEM image at 0 minutes in the top left corner shows the as-deposited seed array at room temperature (30 °C), taken immediately after heating to 90 °C. With the GIS open and stage heated for 10 min, several images of this array were captured every 60 seconds. The images were taken with a field of view of 40 µm x 26.6 µm for a total time of 23.9 s. Thus, 100 pA electron beam current resulted in a dose of 2.2 pC/µm or 14 electrons per nm per single image. Figure SI.3 displays SEM images captured at 3, 7, and 10 mins to visualize the evolution of the autocatalytic gold CVD film growth. The bottom row presents the as-deposited seed array, which was grown on the sample at an elevated temperature of 90 °C, with the same sequence of images taken at different times. These images show that SE imaging acts as iterative seeding to enhance the CVD growth rate significantly compared to the pure autocatalytic growth demonstrated above. Note that when imaging, seeds are also deposited between squares, resulting in a decrease in area-selectivity. By implementing an iterative seeding approach, electrons will only be supplied in the selected areas.
Autocatalytic Control Test
A control test of pattern array deposition was performed without the Au(acac)Me precursor gas being introduced into the vacuum chamber. The array pattern was irradiated with the electron beam while residual chamber gas was present at a base pressure of mbar, both at room temperature (30 °C) and at 90 °C. Subsequently, both irradiated areas were left at elevated temperature of 90 °C for 50 min with the gold precursor gas flow enabled. SE micrographs presenting the results are displayed in Figure SI.4(a) and (b).
Gold particles are observed only for the highest dose as evidenced by the close-up high resolution SE micrograph provided in Figure SI.2(c), which is a magnified view of the highlighted region in Figure SI.2(b). This implies that under high-vacuum (HV) conditions, surface activation by electrons, as also demonstrated by Walz et al. 13, may be achievable. Dark contrast in squares may arise from hydrocarbon contamination on the surface caused by residual gas composition in high vacuum 52. These deposits are commonly referred to as carbonaceous contamination.
Thickness evaluation
Atomic force microscopy (AFM) measurements were carried out to determine the thickness of the autocatalytically grown gold films. An SE image of the autocatalytically grown films at an elevated temperature of 80 °C for 30 min is shown in Figure SI.5(a). Seeding for the films was done at 80 °C. Figure SI.5(b) shows the AFM thickness profiles of the squares in the 3rd and 6th rows, which were seeded with a minimum pitch of 5 nm. The thickness of the film versus the deposition time of the seed layer for each individual square is displayed in Figure SI.5(c). The data was compared to an exponential asymptotic function that exhibits a saturation of thickness. The saturation in autocatalytic CVD gold growth was observed on certain seed squares with a deposition time of 5.6 s for a size of 2 µm x 2 µm, corresponding to an electron dose of 140 pC/µm or about 875 electrons per nm. Corresponding thickness of the film autocatalytically grown for 30 min on this seed layer is about 6.5 nm. Gold films after seeding and autocatalytic growth at 80 °C for longer times of 60 and 90 mins were additionally investigated. A saturation in achievable film thickness in both cases was also observed as discussed in the main manuscript. Autocatalytic growth for 60 min only resulted in additional 0.7 nm, and even 90 minutes total growth time resulted only in 8.5 nm.
Transmission Electron Microscopy
For transmission electron microscopy (TEM) investigations, two different samples were used. Gold films were grown autocatalytically on substrates at a temperature of 90 °C for 30 min, but were seeded at 30 °C for sample (a) and 90 °C for sample (b) (Figure SI.6(a) and (b) respectively). Gold films were grown on a n-type (100) Si chip with native oxide for sample (a) and on a p-type (100) Si chip with native oxide for sample (b). TEM lamellae were prepared according to the standard lift-out protocol using dual beam FIB-SEM Tescan Lyra3. The TEM images were captured using a probe-corrected ThermoFisher Scientific Titan Themis 200 G3. Figure SI.6 shows the full STEM micrographs used for the main manuscript. In both instances, periodic atomic order of specific gold grains can be observed.
Monte-Carlo simulations
The energy deposited by electrons in the material was extracted using a Monte-Carlo method proposed by David C. Joy’s single scattering model 53. The analytical Mott expression 54 implemented the inelastic scattering event, and the modified Bethe expression 53 was used for the elastic energy loss, which is an average energy loss by an electron due to all inelastic scattering events. The structure representing the FEBID-grown pillar was implemented with a cylinder and a truncated cone above the cylinder. Trajectory and energy loss of each electron were recorded and used to calculate the total energy lost by primary electrons on average. Figure SI.7 depicts the paths of 500 electrons with 15 keV energy, traveling within an object that has a top truncated cone with a 10 nm radius and a height of 150 nm, as well as a cylinder with a 25 nm radius. The object has a composition ratio of Au:C as 0.28:0.72, with a density of 6.5 g/cm. The energy transfer value for 4 g/cm lower bound was calculated using the same method. The object’s geometry was derived from SEM images. The energy deposited per each primary electron and line segment between two collisions at the apex of the pillar is shown in the inset of the figure.
Temperature Distribution in Pillars Induced by the Electron Beam
Convection losses in high vacuum ( mbar) are negligible because the mean free path of the residual gas molecules is greater than the size of the vacuum chamber. The radiative losses can be estimated by applying the Stefan-Boltzmann law, where the total radiative power is the difference between the power emitted by the column and the power absorbed by the surrounding medium: , where is the surface area of the column, W/mK is the Stefan-Boltzmann constant, is the emissivity, and is the ambient temperature of the vacuum chamber. The typical pillar deposited at 30 °C has a surface area of m with a radius of nm and a length of nm. In order to estimate the maximum radiative losses, we assume a maximum emissivity of the deposit of 1. For the aforementioned reason, we ignore any irregularities in temperature distribution and assume uniform heating of the entire pillar, with its length being equal to the previously determined temperature of 112 °C (385 K). In this instance, the maximum power output is W. As per Monte-Carlo simulations, a primary electron of 15 kV conveys at least 0.75 kV of its initial energy to the pillar. This then results in an implanted power of W by a focused electron beam with a current of 100 pA. If the power transfer from the focused electron beam is compared with the maximum estimated radiative loss, the latter is at least 2 orders of magnitude lower and thus negligible, which justifies a linear temperature gradient in the pillar.
The following schematics are proposed for a better understanding of the heat transfer and temperature distribution in the electron beam grown pillars. The left-hand side of Figure SI.8(a) depicts the time evolution of a single pillar.
Let us first consider the stage temperature T set at the room temperature of 30 °C. At smaller sizes, the majority of the heat supplied via electron beam is dissipated within the substrate. However, once the pillar reaches a certain height that corresponds to the stopping range of primary electrons in the rod, the implanted energy from primary electrons into the pillar will be maximised. The temperature of the top of the pillar exceeding the base temperature (i.e. the temperature of the substrate) generates an almost linear temperature gradient along the axis of the pillar. This can be attributed to a quasi-continuous heat supply by the electron beam and the stage temperature, T, remaining constant, resulting in a steady solution of the heat transfer equation (cf. 55). This occurs because there are no significant convectional or radiative losses as estimated previously. As a result, the temperature of the upper portion of the pillar may reach the threshold for the autocatalytic growth of gold particles, which only becomes visible once the gold crystals increase to a size that can be observed by SEM imaging at the sidewall of the pillar. During the experiment with Au(acac)Me, this occurred at about 2 min 30 s with a 15 keV 100 pA electron beam. The temperature distribution within the upper section of the pillar remains uncertain because of the presence of gold crystals, which impact heat conductivity.
If a pillar is now deposited with the same deposition parameters and total time, but on the stage at an elevated temperature of 70 °C, an autocatalytic process will take place at a lower point along the pillar axis (see Figure SI.8(a)-(iv)). By progressively increasing the stage temperature, gold crystals will eventually appear on the pillar alongside the sample surface, as shown in Figure SI.8(a)-(v). This enables the calculation of T - the temperature induced by the electron beam at the point along the pillar where the CVD process first occurs (see Figure SI.8(a)). T remains constant during the specified deposition period. If we plot all temperature distributions for each pillar grown at different T (refer to Figure SI.8(b)), we can observe that the point on the pillar heated to the temperature T shifts linearly to the sample surface and collapses with the surface at around 112 °C.